OpenStudy (anonymous):

Find an equation of the line that passes through (2,5) and is parallel to 4x-y=7. Write the equation in slope-intercept form.

OpenStudy (anonymous):

Start by taking 4x-y=7 and isolating y. Once you've done that, what is the slope of that line?

OpenStudy (anonymous):

y=-4x-7/4?

OpenStudy (anonymous):

Close -- first add 'y' to each side, then subtract 7 from each side. What does that give you?

OpenStudy (anonymous):

y=4x+7/4

OpenStudy (anonymous):

It seems like you're dividing 7 by 4, but you shouldn't need to do that -- all you need to do is subtract 7 from each side, not then divide by 4. You're solving for y, not x, so you don't need to divide by 4 to get rid of the "4x".

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

So what do those two equation steps look like when you work them out?

OpenStudy (anonymous):

y-5=4x(x+5) for the point slope right?

OpenStudy (anonymous):

You've got the slope of 4 right. So you can have a basic slope equation of y = 4x. But when you plug in the point (2,5), you're properly subtracting 5 from y. But on the 'x' side, what do you need to do?

OpenStudy (anonymous):

i've got it thanks. 4x-y=7 should be y=4x-7 5=(4)(2)+b 5=8+b -3=b y=4x-3